Find the complete integral of p - 3 x² = q² - y


1
Expert's answer
2022-03-14T17:28:37-0400

The given equation is a separable equation of the form "f(x, p)=g(y, q)".


The complete solution will be, "z = \int pdx + \int qdy + c"


Let "f(x, p)=g(y, q) =a". Solving for "f(x, p)=a~ \&~ g(y, q) =a", we get

"\begin{aligned}
p-3x^2 &= a&; & ~q^2 -y = a\\
p&=3x^2 + a&; & ~q^2=y+a\\
p&=3x^2 + a&; & ~q=\sqrt{y+a}\\ 
\end{aligned}"

The complete solution is,


"\begin{aligned}
z &= \int pdx + \int qdy + c\\
z &= \int (3x^2 + a)dx + \int (\sqrt{y+a})dy + c\\
&= x^{3} + ax + \frac{2}{3}(y+a)^{\frac{3}{2}} + c
\end{aligned}"